OpenStudy (anonymous):

Which Property of Equality should be used first to solve the equation 8 + P divided by 2 = 16 for p? Addition Property of Equality Subtraction Property of Equality Multiplication Property of Equality Division Property of Equality

OpenStudy (anonymous):

@HawkCrimson I THINK I DID IT.

OpenStudy (ikram002p):

\[\frac{ 8+p }{2 }=16 \rightarrow 8+p=32\]

OpenStudy (anonymous):

Simplify the expression -2 + 3(2x – 1) + 6x ?

OpenStudy (anonymous):

it worked! :D first expand the brackets and tell me wat u get :)

OpenStudy (anonymous):

the answer is the multiplication property of equality to get rid of the 2 which is dividing into 8+P

OpenStudy (anonymous):

^ I got that one. Sorry btw I went to go make me food. xD But in my lesson it says to combine like terms. .-.

OpenStudy (anonymous):

so which one are u having problems in?

OpenStudy (anonymous):

Simplify the expression -2 + 3(2x – 1) + 6x ? ^Thats the question I need to do but the stuff theyre telling me to do in the lesson is really confusing so sorry just continue with what you were saying. xD

OpenStudy (anonymous):

so first expand 3(2x - 1) and tell me wat do u get